# The Optimality of Blocking Designs in Equally and Unequally Allocated Randomized Experiments with General Response

### v3

@@ 1.1 Equal Allocation Experimentation
If $n_{T}=n_{C}$, the design is termed equal allocation. The default design here is the balanced complete randomization design (BCRD, Wu, 1981, p. 1171). This design has one constraint: $\sum_{i}w_{i}=0$ which is satisfied by $\binom{2n}{n}$ vectors. These vectors are then sampled uniformly. BCRD has been given the high distinction of “the gold standard”. Cornfield, (1959, p. 245) gives two reasons: (1) known and unknown covariate differences among the groups will be small and bounded and (2) it is the “reasoned basis” for causal inference, a paraphrasing of Fisher, (1935, p. 14). Any other design we will call “restricted” as it is restricted to use fewer assignment vectors.
@@ 2 Problem Setup
We will be analyzing performance for many types of responses. It will be useful to express potential outcome as a sum of two components: (a) what we will call the “response mean” $\mu_{i}$ which is a function of the observed covariates $\boldsymbol{x}_{i}$ and the assignment $w_{i}$ and (b) $z_{i}$ which is a function of unobserved covariates $z_{i}$ and the assignment $w_{i}$. Thus,
$\boldsymbol{y}_{T}=\boldsymbol{\mu}_{T}+\boldsymbol{z}_{T}\in\mathcal{Y},~{}~{}\boldsymbol{y}_{C}=\boldsymbol{\mu}_{C}+\boldsymbol{z}_{C}\in\mathcal{Y}$ (5)
where $\boldsymbol{\mu}_{T}$ and $\boldsymbol{\mu}_{C}$ denote the stacked vectors of the observed covariates’ contribution to the responses when $\boldsymbol{w}=+\boldsymbol{1}_{2n}$ and $\boldsymbol{w}=-\boldsymbol{1}_{2n}$ respectively. Analogously, $\boldsymbol{z}_{T}$ and $\boldsymbol{z}_{C}$ denote the stacked vectors of the unobserved covariates’ contribution to the responses when $\boldsymbol{w}=+\boldsymbol{1}_{2n}$ and $\boldsymbol{w}=-\boldsymbol{1}_{2n}$ respectively. The potential response values belong to different sets $\mathcal{Y}$ depending on response type.
The response types we consider are formally listed in Table 1 (column 1) and this table also lists common generalized linear models (GLMs) for each of the responses’ means (column 4). These means likewise belong to different sets $\Theta$, where $\Theta$ is the space of the $\mu_{i}$’s (column 5). In the population model, the $z_{i}$’s are conceptualized as randomly-realized noise or errors. Regardless of what one believes about their origination, the critical fact is they cannot be modeled with the information the experimenter has at hand. Of course since $z_{i}=y_{i}-\mu_{i}$, the space of permissible values of the $z_{i}$’s is analogously limited for each response type to be within $\mathcal{Y}\,-\,\Theta$ and the space $\mathcal{Y}$ will be different for each response type.
@@ 3 Analyzing Designs
It is impossible to employ the MSE as the criterion in which we compare different designs since $\boldsymbol{z}_{T}$ and $\boldsymbol{z}_{C}$ are not observed. Thus, as in our previous work, we do not employ MSE as the performance criterion directly but instead remove the dependence on the $z_{i}$’s in three different ways to arrive at three different practical criterions (a) the worst-case MSE, (b) the mean MSE and (c) the tail MSE. These are explored in the next three sections respectively.
@@ 3.1 The Worst Case MSE
which is exactly tractable for the case of the response being continuous. In this case, the inner maximum is the largest eigenvalue of $\boldsymbol{\Sigma}_{W}$. The $\boldsymbol{W}$ that minimizes this largest eigenvalue is such where each subject is assigned on the basis of a fair coin flip and is either called the complete randomization design (Rosenberger and Lachin, , 2016, Chapter 3.2) or the the Bernoulli Trial (Imbens and Rubin, , 2015, Chapter 4.2). In this design, $\boldsymbol{W}=\boldsymbol{I}_{2n}$ which mirrors the minimax result found in Kallus, (2018, Section 2.1). However, this design violates our space of considered designs which are only those that satisfy a prespecified allocation to the two treatments (Assumption 2.1). Within the space of allowable designs, we have the following no free lunch result which follows from Kapelner et al., (2021, Theorem 2.1) and proven in Section A.3 in the Supplemental Material.
@@ 3.2 The Mean MSE
where $\boldsymbol{\mu}:=\boldsymbol{\mu}_{T}/r+\boldsymbol{\mu}_{C}/\tilde{r}$ and $c_{Z}$ is a constant with respect to the design $\boldsymbol{W}$ defined in Section A.5. Intuitively, this expectation operation renders the third term in the MSE decomposition of Equation 7 as a constant and eliminates the second term. Thus, the optimization problem can then be expressed as $\operatorname*{arg\,min~{}}_{\boldsymbol{W}\in\mathcal{W}}\boldsymbol{\mu}^{\top}\boldsymbol{\Sigma}_{W}\boldsymbol{\mu}$. Since the second term in $\boldsymbol{\Sigma}_{W}$ is a constant with respect to the design and the first term can be written as in Equation 6, this optimization problem can alternatively be expressed as
$\boldsymbol{W}_{*}=\operatorname*{arg\,min~{}}_{\boldsymbol{W}\in\mathcal{W}}\boldsymbol{\mu}^{\top}\mathbb{E}\left[\boldsymbol{W}\boldsymbol{W}^{\top}\right]\boldsymbol{\mu}=\operatorname*{arg\,min~{}}_{\boldsymbol{W}\in\mathcal{W}}\sum_{k=1}^{N_{\boldsymbol{W}}}p_{k}(\boldsymbol{w}_{k}^{\top}\boldsymbol{\mu})^{2}=\operatorname*{arg\,min~{}}_{\boldsymbol{w}~{}:~{}\boldsymbol{w}^{\top}\boldsymbol{1}_{2n}=(r-\tilde{r})n}(\boldsymbol{w}^{\top}\boldsymbol{\mu})^{2}.$ (16)
where the sum in the middle step above is taken over all allocations in the design and the $p_{k}$’s denote the probability of the allocation being selected in the experiment.
@@
$\boldsymbol{\mu}=\left((\beta_{0}+\beta_{T})/r+(\beta_{0}-\beta_{T})/\tilde{r}\right)\boldsymbol{1}_{2n}+\beta_{1}(1/r+1/\tilde{r})\boldsymbol{x}.$
@@ 3.3 The MSE’s Tail
$\text{Quantile}_{\boldsymbol{Z}_{T},\boldsymbol{Z}_{C}}\left[\mathbb{M}\text{SE}_{\boldsymbol{W}}\left[\hat{\tau}\right],q\right]=\mathbb{E}_{\,\boldsymbol{Z}_{T},\boldsymbol{Z}_{C}}\left[\mathbb{M}\text{SE}_{\boldsymbol{W}}\left[\hat{\tau}\right]\right]+c_{\boldsymbol{W},q}\sqrt{\mathbb{V}\text{ar}_{\boldsymbol{Z}_{T},\boldsymbol{Z}_{C}}\left[\mathbb{M}\text{SE}_{\boldsymbol{W}}\left[\hat{\tau}\right]\right]}.$ (19)
There exists a constant $c_{\boldsymbol{W},q}$ that depends upon the design and the quantile desired. As this constant is not tractable to analyze, we consider the criterion we call the approximate tail criterion where we view this constant as fixed with respect to the design,
@@
This criterion allows us to incorporate the effect of both the mean and variance with respect to the unobserved covariates and allows us to compare designs asymptotically. We discuss other criterions that can capture both considerations in Section 5. We demonstrate later via simulation that employing the standard normal quantile $c_{q}:=-z_{1-q}$ provides near-exact quantiles as we conjecture there is a central limit theorem on $\mathbb{M}\text{SE}_{\boldsymbol{W}}\left[\hat{\tau}\right]$ over realizations of $\boldsymbol{Z}_{T},\boldsymbol{Z}_{C}$ if $N_{\boldsymbol{W}}$ is large.
@@ 3.3.5 The Asymptotically Optimal Block Design
The notation $a_{n}\approx b_{n}$ denotes that the limit of $a_{n}/b_{n}$ is finite and positive. Assumption 3.4 connects the number of blocks in the design and the balance term $\mathcal{B}_{1}$. As shown below, it holds if the vector $\boldsymbol{\mu}$ is approximately equally spaced, meaning that $\mu_{j}\approx M/B$ (recall that we assumed in Equation 17 that $\boldsymbol{\mu}$ is ordered). We discuss below a relaxation of this assumption. Assumption 3.4 is motivated by noticing that
@@
where $\boldsymbol{\mu}_{b}$ is the sub-vector of $\boldsymbol{\mu}$ that corresponds to the $b$-th block is reasonable. This is because for $b=1$ (and similarly for the other block sizes) $\mu_{1,1},\ldots,\mu_{1,n_{B}}$ are in $[0,\mu_{n_{B}}]$ and $\mu_{n_{B}}\approx M/B$ if the $\mu$’s are ordered approximately uniformly in $[0,M]$. Therefore, the sample variance satisfies
@@
Assumption 3.4 is reasonable when the $\mu_{i}$’s are ordered as in Equation 17 and are approximately uniform. If this is not the case, it is expected that $\mathcal{B}_{1}$ decreases slower with $B$, and then the optimal block size might be of larger order than $n^{1/4}$. For example, if $\mathcal{B}_{1}\approx n/B$ (instead of $n/B^{2}$), then similar arguments used to prove Theorem 3.6 will instead imply that the optimal block size $B$ satisfies $B/n^{1/2}\to\infty$ and $B/n\to 0$.
We discuss now violations of our assumed knowledge of the order of the $\mu_{i}$’s (Equation 17). The results above imply if the $\mu_{i}$’s are neither ordered nor approximately uniform, in order for block designs to perform optimally, it is enough to require that there exists a block design with $B/n\to 0$ such that its balance term $\mathcal{B}_{1}$ is $o(\sqrt{n})$ as summarized in Remark 3.1 below.
@@ 4.2 Block design construction for $p=1$
Simply ordering the subjects by the order of the one covariate measurement then splicing into $B$ even sets provides the block design structure as all observed covariate response functions are monotonic in $\boldsymbol{x}$ (see columns 3 and 4 of Table 1). Thus, the entries of $\boldsymbol{\mu}$ are ordered within the blocks corresponding to the block-diagonal structure of $\boldsymbol{\Sigma}_{W}$. This order is required for our theoretical results to hold. Simulation results for $p=1$ constitute empirical confirmation of our mathematical results.
@@ 4.4 Results
For $p>1$, our theorems do not precisely apply as the $\mu_{i}$’s are not precisely ordered into blocks, but only approximately by the arrangement of blocks of two covariates. Notwithstanding, the results for $p>1$ overall demonstrate the robustness of our theoretical results to situations of “imperfect” blocking; and imperfect blocking is the design employed in all real-world experimental settings. Thus, the $p>1$ results are an empirical demonstration of Remark 3.1 which provides evidence that our theory applies to real-world block designs that imperfectly order $\boldsymbol{\mu}$.
@@
For both equal and unequal allocation, the approximate tail criterion computed using the sample average, sample standard deviation and $c_{q}=1.645$ (the standard normal 95% quantile) is nearly the same as the empirical 95%ile (compare red and blue lines for all plot windows in Figures 2(a) and 2(b)). This comparison demonstrates that the approximate tail criterion (Equation 20) is nearly equal to the actual quantile tail criterion (Equation 19) for the gamut of real world scenarios in all response types. Whatever difference exists between the approximate quantile and the actual quantile further does not affect the optimal lowest MSE designs (visualized as the $B$ value corresponding to the minimum of the curves).
@@ 5 Discussion
Herein, we derive the quantile tail criterion for general response types. We then prove that the asymptotically optimal design is achieved with a block design with very few blocks — on the order of between $n^{1/4}$ and $n$ as seen in case (iii) of Theorem 3.6. Blocking with few blocks is an intuitive imbalance-reducing idea proposed by Fisher nearly 100 years ago and is one of the most popular designs employed today in clinical trials. This work vindicates this common practice for all common response types by analyzing its performance from a new angle. We demonstrate that blocking is the optimal design harmonizing the degree of randomness and imbalance reduction. Although this work focuses on estimation, these blocking designs should have high power as well as many allocations can be generated which are highly independent (Krieger et al., , 2020).
@@
There are many other avenues to explore. First, although we proved that blocking with few blocks are asymptotically optimal, we did not prove it is uniquely optimal. Other designs may also satisfy the lower bound of Equation 23 and offer better finite sample properties. Other optimal designs are additionally of profound importance because blocking suffers from two main practical problems (1) the curse of dimensionality as it can only handle $p\leq 3$ in practice and (2) the design requires arbitrary cut-points if the covariate is continuous. Further, our definition of the blocking design has strong assumptions (Assumptions 3.4 and 3.5) that are needed to prove its asymptotic optimality. These assumptions appear to be ostensibly realistic as demonstrated by the simulations. However, we can explore proving the results under weaker assumptions. Another area of exploration could be understanding the tail criterion for other designs such as rerandomization and greedy pair switching. Additionally, survival is an important experimental endpoint in clinical trials but this work only considered uncensored survival measurement while the vast majority of survivals collected in real-world experiments are censored. Also, we did not investigate the distribution of the $\mu_{i}$’s relative to the optimal block size and we believe this can be a source of fruitful future work. In our previous work (Section 2.3), we explored the ordinary least squares estimator. This avenue would be difficult to explore for general response types as the analogous estimators do not have closed form (e.g. if the response is binary, the multivariate logistic regression estimates are the result of computational iteration). It is our intuition the theorems found herein would extend to these commonly used covariate-controlled estimators, but more work is necessary. Further, our tail criterion is only one metric that combines mean and standard deviation of MSE. Other criterions can be explored e.g. a tail simultaneous in allocation and unmeasured covariates, the ratio of mean to standard deviation (the coefficient of variation), etc. Also, work can be done to find the optimal imbalance design under unequal allocation, the analogue of PB in equal allocation (see discussion below Equation 16). Finally, in the case of unequal allocation, we can develop a PM analogue by performing $(n_{C}/n_{T}):1$ matching. If $n_{C}/n_{T}$ is not a natural number, some matches will have $\lfloor n_{C}/n_{T}\rfloor$ control subjects for one treatment subject and some matches can have $\lceil n_{C}/n_{T}\rceil$ control subjects for one treatment subject (in approximately the correct proportion). An algorithm to create these tuplets based on intra-subject observed covariate distance is only known to be able to be solved in polynomial time when each matched set has one treated and one control. Clearly, this more general $(n_{C}/n_{T}):1$ matching problem is more complex and might very well be NP-hard.

### v4

@@ 1.1 Equal Allocation Experimentation
If $n_{T}=n_{C}$, the design is termed equal allocation. The default design here is the balanced complete randomization design (BCRD, (Wu, , 1981, p. 1171)). This design has one constraint: $\sum_{i}w_{i}=0$ which is satisfied by $\binom{2n}{n}$ vectors. These vectors are then sampled uniformly. BCRD has been given the high distinction of “the gold standard”. Cornfield, (1959, p. 245) gives two reasons: (1) known and unknown covariate differences among the groups will be small and bounded and (2) it is the “reasoned basis” for causal inference, a paraphrasing of Fisher, (1935, p. 14). Any other design (besides the Bernoulli Trial design discussed in Section 3.1) we will call “restricted” as it is restricted to use fewer assignment vectors.
@@ 2 Problem Setup
We will be analyzing performance for many types of responses. It will be useful to express potential outcome as a sum of two components: (a) $\mu_{T,i},\mu_{C,i}$ which are a function of the observed covariates $\boldsymbol{x}_{i}$ (b) $z_{T,i},z_{C,i}$ that which is a function of the unobserved covariates. Letting $\boldsymbol{\mu}_{T}$ and $\boldsymbol{\mu}_{C}$ (and $\boldsymbol{z}_{T}$ and $\boldsymbol{z}_{C}$) denote the stacked vectors of the observed (and unobserved) covariates’ contribution to the responses when $\boldsymbol{w}=+\boldsymbol{1}_{2n}$ and $\boldsymbol{w}=-\boldsymbol{1}_{2n}$ respectively, we have
$\boldsymbol{y}_{T}=\boldsymbol{\mu}_{T}+\boldsymbol{z}_{T}\in\mathcal{Y},~{}~{}\boldsymbol{y}_{C}=\boldsymbol{\mu}_{C}+\boldsymbol{z}_{C}\in\mathcal{Y}.$ (5)
The response types we consider are formally listed in Table 1 (column 1). The potential response values belong to different sets $\mathcal{Y}$ depending on response type (column 2). This table also lists common generalized linear models (GLMs) for each of the responses’ means (column 4). These means likewise belong to different sets $\Theta$, where $\Theta$ is the space of the $\mu_{T,i},\mu_{C,i}$’s (column 5). In the population model, the $z_{T,i},z_{C,i}$’s are conceptualized as randomly-realized noise or errors. Regardless of what one believes about their origination, the critical fact is they cannot be modeled with the information the experimenter has at hand.
@@ 3 Analyzing Designs
It is impossible to employ the MSE as the criterion in which we compare different designs since $\boldsymbol{z}_{T}$ and $\boldsymbol{z}_{C}$ are not observed. Thus, as in our previous work, we do not employ MSE as the performance criterion directly but instead remove the dependence on $\boldsymbol{z}_{T},\boldsymbol{z}_{C}$ in three different ways to arrive at three different practical criterions (a) the worst-case MSE, (b) the mean MSE and (c) the tail MSE. These are explored in the next three sections respectively.
@@ 3.1 The Worst Case MSE
which is exactly tractable for the case of the response being continuous. In this case, the inner maximum is the largest eigenvalue of $\boldsymbol{\Sigma}_{W}$. The $\boldsymbol{W}$ that minimizes this largest eigenvalue is such where each subject is assigned on the basis of a fair coin flip and is either called the complete randomization design (Rosenberger and Lachin, , 2016, Chapter 3.2) or the the Bernoulli Trial (Imbens and Rubin, , 2015, Chapter 4.2). In this design, $\boldsymbol{\Sigma}_{W}=\boldsymbol{I}_{2n}$ which mirrors the minimax result found in Kallus, (2018, Section 2.1). However, this design violates our space of considered designs which are only those that satisfy a prespecified allocation to the two treatments (Assumption 2.1). Within the space of allowable designs, we have the following no free lunch result which follows from Kapelner et al., (2021, Theorem 2.1) and proven in Section A.3 in the Supplemental Material.
@@ 3.2 The Mean MSE
where $\boldsymbol{\mu}:=\boldsymbol{\mu}_{T}/r+\boldsymbol{\mu}_{C}/\tilde{r}$ and $c_{Z}$ is a constant with respect to the design $\boldsymbol{W}$ defined in Section A.5. Intuitively, this expectation operation renders the third term in the MSE decomposition of Equation 7 as a constant and eliminates the second term. Thus, the optimization problem can then be expressed as $\operatorname*{arg\,min~{}}_{\boldsymbol{W}\in\mathcal{W}}\boldsymbol{\mu}^{\top}\boldsymbol{\Sigma}_{W}\boldsymbol{\mu}$. Since the second term in $\boldsymbol{\Sigma}_{W}$ is a constant with respect to the design and the first term can be written as in Equation 6, this optimization problem can alternatively be expressed as the design that minimizes
$\boldsymbol{\mu}^{\top}\mathbb{E}\left[\boldsymbol{W}\boldsymbol{W}^{\top}\right]\boldsymbol{\mu}=\sum_{k=1}^{N_{\boldsymbol{W}}}p_{k}(\boldsymbol{w}_{k}^{\top}\boldsymbol{\mu})^{2}.$ (16)
Thus this design chooses, with equal probability, $\boldsymbol{w}$’s from the set
$\operatorname*{arg\,min~{}}_{\left\{\boldsymbol{w}~{}:~{}\boldsymbol{w}^{\top}\boldsymbol{1}_{2n}=(r-\tilde{r})n\right\}}(\boldsymbol{w}^{\top}\boldsymbol{\mu})^{2}.$
@@
$\boldsymbol{\mu}=\left(\frac{\beta_{0}+\beta_{T}}{r}+\frac{\beta_{0}-\beta_{T}}{\tilde{r}}\right)\boldsymbol{1}_{2n}+\beta_{1}\left(\frac{1}{r}+\frac{1}{\tilde{r}}\right)\boldsymbol{x}.$
@@ 3.3 The MSE’s Tail
$\text{Quantile}_{\boldsymbol{Z}_{T},\boldsymbol{Z}_{C}}\left[\mathbb{M}\text{SE}_{\boldsymbol{W}}\left[\hat{\tau}\right],q\right]=\mathbb{E}_{\,\boldsymbol{Z}_{T},\boldsymbol{Z}_{C}}\left[\mathbb{M}\text{SE}_{\boldsymbol{W}}\left[\hat{\tau}\right]\right]+c_{\boldsymbol{W},q}\sqrt{\mathbb{V}\text{ar}_{\boldsymbol{Z}_{T},\boldsymbol{Z}_{C}}\left[\mathbb{M}\text{SE}_{\boldsymbol{W}}\left[\hat{\tau}\right]\right]}$ (19)
where $c_{\boldsymbol{W},q}$ is a constant that depends upon the design and the quantile desired. As this constant is not tractable to analyze, we consider the criterion we call the approximate tail criterion where we view this constant as fixed with respect to the design,
@@
This criterion allows us to incorporate the effect of both the mean and variance with respect to the unobserved covariates and allows us to compare designs asymptotically. We discuss why this approximation is compelling along with alternative criterions that capture both considerations in Section 5.
@@ 3.3.5 The Asymptotically Optimal Block Design
The notation $a_{n}\approx b_{n}$ denotes that the limit of $a_{n}/b_{n}$ is finite and positive. Assumption 3.4 connects the number of blocks in the design and the balance term $\mathcal{B}_{1}$. As shown below, it holds if the vector $\boldsymbol{\mu}$ is approximately equally spaced, meaning that $\mu_{j}\approx M/B$ (which is reasonable under the assumption that that $\boldsymbol{\mu}$ is ordered; see Equation 17). Assumption 3.4 is motivated by noticing that
@@
where $\boldsymbol{\mu}_{b}$ is the sub-vector of $\boldsymbol{\mu}$ that corresponds to the $b$-th block is a reasonable assumption. This is because for $b=1$ (and similarly for the other block sizes) $\mu_{1,1},\ldots,\mu_{1,n_{B}}$ are in $[0,\mu_{n_{B}}]$ and $\mu_{n_{B}}\approx M/B$ if the $\mu$’s are ordered approximately uniformly in $[0,M]$. Therefore, the sample variance satisfies
@@
Assumption 3.4 is reasonable when the $\mu_{i}$’s are ordered as in Equation 17 and are approximately uniform. If this is not the case, it is expected that $\mathcal{B}_{1}$ decreases slower with $B$, and then the optimal block size might be of larger order than $n^{1/4}$. For example, if $\mathcal{B}_{1}\approx n/B$ (instead of $n/B^{2}$), then similar arguments used to prove Theorem 3.6 will instead imply that the optimal block size $B$ satisfies $B/n^{1/2}\to\infty$ and $B/n\to 0$. More generally, the results above imply if the $\mu_{i}$’s are neither ordered nor approximately uniform, in order for block designs to perform optimally, it is enough to require that there exists a block design with $B/n\to 0$ such that its balance term $\mathcal{B}_{1}$ is $o(\sqrt{n})$ as summarized in Remark 3.1 below.
@@ 4.2 Block design construction for $p=1$
Simply ordering the subjects by the order of the one covariate measurement then splicing into $B$ even sets provides the block design structure as all observed covariate response functions are monotonic in $\boldsymbol{x}$ (see columns 3 and 4 of Table 1). Thus, the entries of $\boldsymbol{\mu}$ are ordered within the blocks corresponding to the block-diagonal structure of $\boldsymbol{\Sigma}_{W}$. Simulation results for $p=1$ constitute empirical confirmation of our mathematical results.
@@ 4.4 Results
For $p>1$, the $\mu_{i}$’s are not precisely ordered into blocks, but only approximately by the arrangement of blocks of two covariates. Notwithstanding, the results for $p>1$ overall demonstrate the robustness of our theoretical results to situations of “imperfect” blocking; and imperfect blocking is the design employed in all real-world experimental settings. Thus, the $p>1$ results are an empirical demonstration of Remark 3.1 which provides evidence that our theory applies to real-world block designs that imperfectly order $\boldsymbol{\mu}$.
@@
For both equal and unequal allocation, the approximate tail criterion computed using the sample average, sample standard deviation and $c_{q}=1.645$ (the standard normal 95% quantile) is nearly the same as the empirical 95%ile (compare red and blue lines for all plot windows in Figures 1(a) and 1(b)). This comparison demonstrates that the approximate tail criterion (Equation 20) is nearly equal to the actual quantile tail criterion (Equation 19) for the gamut of real world scenarios in all response types. Whatever difference exists between the approximate quantile and the actual quantile further does not affect the optimal lowest MSE designs (visualized as the $B$ value corresponding to the minimum of the curves).
@@ 5 Discussion
Herein, we derive an approximate quantile tail criterion for general response types. We then prove that the asymptotically optimal design is achieved with a block design with very few blocks — on the order of between $n^{1/4}$ and $n$ as seen in case (iii) of Theorem 3.6. Blocking with few blocks is an intuitive imbalance-reducing idea proposed by Fisher nearly 100 years ago and is one of the most popular designs employed today in clinical trials. This work vindicates this common practice for all common response types by analyzing its performance from a new angle. We demonstrate that blocking is the optimal design harmonizing the degree of randomness and imbalance reduction. Although this work focuses on estimation, these blocking designs should have high power as well as many allocations can be generated which are highly independent (Krieger et al., , 2020).
[…]
Regardless, we believe such pathological settings will be rare as theoretically discussed in Sharpe, (1970) and as empirically demonstrated by the simulations of Section 4. Here, we found that employing $c_{q}$ as the standard normal quantile provides near-exact quantiles across the block designs considered. We conjecture there is a central limit theorem on $\mathbb{M}\text{SE}_{\boldsymbol{W}}\left[\hat{\tau}\right]$ over realizations of $\boldsymbol{Z}_{T},\boldsymbol{Z}_{C}$ if $N_{\boldsymbol{W}}$ is large; this can be left to future work.
In Section 3.2 we assumed knowledge of the order $\boldsymbol{\mu}$, which is reasonable when $p=1$ but not generally. When this assumption holds and $\boldsymbol{\mu}$ is approximately equally spaced (see Assumption 3.4) then Theorem 3.6 implies that the optimal number of blocks is between $n^{1/4}$ and $n$. More generally, Remark 3.1 shows that blocking designs with number of blocks of smaller order than $n$ and that are balanced enough are asymptotically optimal (in the sense that they asymptotically achieve the lower bound of Theorem 3.4). The results of Section 4 demonstrate that in the models we considered, indeed blocking design with only a few number of blocks (e.g., 4 or 8) are optimal (among blocking designs) also for the case $p>1$.
@@
There are many other avenues to explore. First, although we proved that blocking with few blocks are asymptotically optimal, we did not prove it is uniquely optimal. Other designs may also satisfy the lower bound of Equation 23 and offer better finite sample properties. Other optimal designs are additionally of profound importance because blocking suffers from two main practical problems (1) the curse of dimensionality as it can only handle $p\leq 3$ in practice and (2) the design requires arbitrary cut-points if the covariate is continuous. Another area of exploration could be understanding the tail criterion for other designs such as rerandomization and greedy pair switching. Additionally, survival is an important experimental endpoint in clinical trials but this work only considered uncensored survival measurement while the vast majority of survivals collected in real-world experiments are censored. Also, we did not investigate the distribution of the $\mu_{i}$’s relative to the optimal block size and we believe this can be a source of fruitful future work. In our previous work (Section 2.3), we explored the ordinary least squares estimator. This avenue would be difficult to explore for general response types as the analogous estimators do not have closed form (e.g. if the response is binary, the multivariate logistic regression estimates are the result of computational iteration). It is our intuition the theorems found herein would extend to these commonly used covariate-controlled estimators, but more work is necessary. Further, our tail criterion is only one metric that combines mean and standard deviation of MSE. Other criterions can be explored e.g. a tail simultaneous in allocation and unmeasured covariates, the ratio of mean to standard deviation (the coefficient of variation), etc. Also, work can be done to find the optimal imbalance design under unequal allocation, the analogue of PB in equal allocation (see discussion below Equation 16). Finally, in the case of unequal allocation, we can develop a PM analogue by performing $(n_{C}/n_{T}):1$ matching. If $n_{C}/n_{T}$ is not a natural number, some matches will have $\lfloor n_{C}/n_{T}\rfloor$ control subjects for one treatment subject and some matches can have $\lceil n_{C}/n_{T}\rceil$ control subjects for one treatment subject (in approximately the correct proportion). An algorithm to create these tuplets based on intra-subject observed covariate distance is only known to be able to be solved in polynomial time when each matched set has one treated and one control. Clearly, this more general $(n_{C}/n_{T}):1$ matching problem is more complex and might very well be NP-hard.